# Rise and Fall of Anderson Localization By Lattice Vibrations: A Time-Dependent Machine Learning Approach

### v1

@@ Abstract
The intricate relationship between electrons and the crystal lattice is a linchpin in condensed matter, traditionally described by the Fröhlich model encompassing the lowest-order lattice-electron coupling. Recently developed quantum acoustics emphasizing the wave nature of lattice vibrations has enabled the exploration of previously uncharted territories of electron-lattice interaction, not accessible with conventional tools such as perturbation theory. In this context, our agenda here is two-fold. First, we showcase the application of machine learning methods to categorize various interaction regimes within the subtle interplay of electrons and the dynamical lattice landscape. Second, we shed light on a nebulous region of electron dynamics identified by the machine learning approach and then attribute it to transient localization, where strong lattice vibrations result in a momentary Anderson prison for electronic wavepackets, later released by the evolution of the lattice. Overall, our research illuminates the spectrum of dynamics within the Fröhlich model, such as transient localization, which has been suggested as a pivotal factor contributing to the mysteries surrounding strange metals. Furthermore, this paves the way for utilizing time-dependent perspectives in machine learning for designing materials with tailored electron-lattice properties.
@@ I Introduction
Anderson localization refers to the cessation of diffusive wave propagation in disordered systems. [1] However, achieving Anderson localization requires the potential to be stationary and the absence of interactions, e.g. between particles or time-dependent external fields. In reality, electrons interact with lattice vibrations, making the localization of electrons in solids challenging. The balance between being localized or not was recognized early on by researchers like Gogolin [2, 3], Thouless [4], and even Anderson himself [1, 5]. Nowadays, the complex interplay between Anderson localization and lattice vibrations is observed in various random metal alloys and other disordered systems, such as crystalline organic semiconductors [6, 7] and halide perovskites [8]. The random fluctuations caused by lattice motion gradually disrupt the quantum interference necessary for electronic state localization, leading to what has been coined transient localization (for capturing the essential aspects, see, e.g., Ref. [9]). This phenomenon combines aspects of both Anderson localized and itinerant electron systems.
@@
While the Anderson localization is typically explored within the framework of a tight-binding scheme featuring random on-site energies, the standard model for lattice vibrations is established by Fröhlich, which encompasses the linear coupling between an electron and the lattice. Conventionally, lattice vibrations are viewed through a number state perspective, but the coherent state representation introduced in Ref. [10], known as quantum acoustics, treats lattice vibrations as waves rather than individual phonons. This perspective unveils a duality between particle and wave pictures akin to quantum optics established by Glauber. Moreover, it allows for the electron-lattice interactions to be described in terms of a quasi-classical internal field, reminiscent of Bardeen and Shockley’s concept regarding dynamical lattice distortions in nonpolar semiconductors [11, 12]. In particular, the deformation potential arising from lattice vibrations enables a quantum-coherent, nonperturbative treatment of charge carriers in coordinate space. In addition to recovering the results of the conventional Bloch-Grüneisen thory [10], the program of quantum acoustics has illuminated mysteries surrounding strange metals where transient localization plays a central role, such as T-linear resistivity at the Planckian limit surpassing the Mott-Ioffe-Regel threshold [13], and a shift of the Drude peak in the optical conductivity towards the infrared range [14]. Motivated by these advancements, we aim to identify various classes of dynamics hidden within the eminent Fröhlich model we express in the coherent state representation.
@@
The quantum acoustical approach enables the generation of large amounts of time-dependent charge carrier wavefunctions as a function of the system parameters. Clustering, which is a common unsupervised learning technique, provides an effective means to explore the spectrum of carrier behavior by grouping similar dynamical profiles into clusters. In general, unsupervised machine learning (ML) methods have been established as a powerful tool to identify complex patterns in large unstructured data sets [15, 16, 17].
@@
In the broader landscape of ML applications in physics, our approach aligns with the recent uses of machine learning to understand and categorize complex physical phenomena, such as many-body localization and phase transitions [17, 18, 19, 20, 21, 22]. However, it is important to distinguish our work from the common narrative of “using ML to do physics.” Instead, our method uses ML as a tool that complements traditional analytical and numerical methods. This distinction underscores a shift from merely applying ML techniques to physics problems towards a more integrated approach where ML assists in how we conceptualize and explore physical systems. We would also like to note that this study represents one of the first applications of ML techniques to the analysis of condensed matter systems through a time-dependent lens. Moreover, our approach not only goes beyond the established focus on eigenstates but also extends the application of these techniques to many-body systems outside of tight-binding models, such as spin chains.
@@
Our program is as follows. In Sec. II, we delineate the theoretical framework across three stages. We first put forward the concept of deformation potential (Sec. II.1), subsequently highlighting its significance as a palpable non-perturbative internal field for electrons (Sec. II.2). To facilitate the analysis of electron-lattice dynamics, we introduce a machine learning methodology in Sec. II.3. In Sec. III, we present our classification of wavepacket dynamics leveraging the ML approach, exploring variations in the strength of the electron-lattice interaction and illustrating a resulting “phase diagram”. Additionally, we conduct a detailed examination of one of the identified sectors connected to transient localization. Finally, we conclude our findings and discussions in Sec. IV.
@@ II Theory and Methods
where $c_{\mathbf{p}}$ $(c_{\mathbf{p}}^{\dagger})$ is the creation (annihilation) operator for electrons with momentum $\mathbf{p}$ and energy $\varepsilon_{\mathbf{p}}$; whereas $a_{\mathbf{q}}$ $(a_{\mathbf{q}}^{\dagger})$ is the creation (annihilation) operator for longitudinal acoustic phonons of wave vector $\mathbf{q}$ and energy $\hbar\omega_{\mathbf{q}}$, respectively. The electron-phonon interaction is defined by its Fourier components $g_{\mathbf{q}}$. This Hamiltonian embodies the lattice $\mathbf{q}$, the electrons $\mathbf{p}$, and their lowest-order (linear) interaction that we next cast into the multimode coherent state basis of lattice degrees of freedom $|\chi\rangle$.
@@ II.1 Deformation Potential
The coherent state picture developed in Ref. [10] is the dual partner of the traditional number state description of electron-lattice dynamics. In this framework, each normal mode of lattice vibration with a wave vector $\mathbf{q}$ is associated with a coherent state $|\mathbf{q}\rangle$. At thermal equilibrium, each mode can be considered to be equilibrated with a heat bath at temperature $T$, giving thermal ensembles of coherent states where the average occupation of the mode $\langle n_{\mathbf{q}}\rangle_{\textrm{th}}$ is given by the Bose-Einstein distribution. Employing the independence of normal modes, entire lattice vibrations can be described as the product state of the coherent states of the normal modes, in other words as a multimode coherent state $|\chi\rangle=\bigotimes_{\mathbf{q}}|\mathbf{q}\rangle$, as studied in Ref [23].
Even though the Fock state perspective focusing on the particle characteristics of lattice vibrations and the coherent state viewpoint emphasizing the wave nature are formally equivalent, the approximations that they inspire are vastly different. As elaborated in Ref. [10], the quasi-classical limit of quantum acoustics unveils a real-space, time-dependent description of electron-lattice interaction in terms of the deformation potential
@@
where $\varphi_{\mathbf{q}}$ is the phase of the coherent state $|\mathbf{q}\rangle$. We furthermore assume the phases $\varphi_{\mathbf{q}}$ to be uniformly distributed random variables, and employ the Debye model, assuming the linear dispersion $\omega_{\mathbf{q}}=v_{s}|\mathbf{q}|$, where $v_{s}$ is the speed of sound. Therefore, the time dependence of the deformation potential is governed by the wave equation:
@@
Figure 1: Snapshots of the deformation potential at three different example temperatures. The left and middle panels demonstrate the awakening of new vibrational modes with increasing temperature, giving rise to finer details in the potential. At the same time, the bumps and dips of the potential become higher and deeper. On the other hand, when the Debye temperature is reached, all the modes are active, and the potential features grow as $\sim\sqrt{T}$, as shown in the right panel. For the sake of illustration, the left panel has a different color scale than the middle and right panels.
@@
The deformation potential in itself is a peculiar object. For instance, it is homogeneously random, meaning that the probability distribution of potential values $V_{D}$ does not depend on a position $\mathbf{r}$ or time $t$ (with the assumption of random phases). Therefore, each spatio-temporal patch of the potential is statistically indistinguishable from another. The typical length scale of the spatial correlations is determined by its largest wavenumber components $\sim q_{D}$. Similarly, the typical timescale of the potential change is determined by its largest frequency components $\sim\omega_{D}$. This special type of spatial-temporal correlation sets the deformation potential apart from other types of lattice distortions, as commonly investigated in the context of Anderson localization [1].
@@ II.2 Electron Dynamics
where $m^{*}$ is the effective (band) mass of the electron and $V_{D}(\mathbf{r},t)$ the deformation potential given by Eq. (2). This effective Hamiltonian $\mathcal{H}_{0}$ represents the electron component of the Fröhlich model defined previously in Eq. (1) within the framework of the effective mass approximation.
Figure 2: The quantum acoustical Wave-on-Wave (WoW) approach to charge carrier dynamics. An electron wavepacket propagates atop a deformation potential, which itself evolves according to the wave equation. As it traverses this shifting acoustic landscape, shaped by acoustic deformations, the electron undergoes quasi-elastic scattering akin to impurity scattering.
Our investigation of electron dynamics under the defined effective Hamilton $\mathcal{H}_{0}$ approaches the issue from the point of view of Gaussian wavepackets that are a common tool for analyzing time-dependent aspects of a quantum system [27, 28], for instance in the studies of quantum optics [29, 30], scarring [31, 32, 33] and branched flow [34, 35, 36]. Here, we choose the following test Gaussian for representing the charge carrier:
@@
where $\mathcal{N}$ is the normalization factor, $\bm{\sigma}=(\sigma_{x}^{-1},\sigma_{y}^{-1})$ describing the initial width of the wavepacket. Without loss of generality, we can choose to launch the test wavepacket into the x direction with the Fermi momentum, thus $\mathbf{k}=(k_{F},0)$, where $k_{F}$ is the Fermi wavevector. The memory of the initial form of the wavepacket is quickly lost in the chaotic potential and its exact form is unimportant.
@@
Figure 2 illustrates the charge carrier wave, originally a Gaussian, described in Eq. (4) evolving under the influence of the dynamic lattice wave field, converting the always-accessible wave nature of lattice vibrations into something valuable, a point where quantum-acoustical perspective becomes tangible. Within this kind of Wave-on-Wave (WoW) approach, as detailed in Refs. [14, 13], one winds up solving two interacting equations of motions simultaneously, one for the lattice and one for the electron. The dynamic deformation potential is separately solved by employing a finite difference wave equation solution. Two coupled equations (i.e. time-dependent Schrödinger equation for electrons, wave equation for the lattice vibrations) are solved independently but simultaneously with mean field back-action to model nonperturbatively the coupled coherent electron-lattice dynamics.
@@
On the other hand, the comparison of the kinetic energy of the electron (a fair approximation given by the Fermi energy) with the root mean square of the deformation potential determines whether the lattice vibration and its resultant electron scattering can be treated perturbatively or not. In essence, the deformation potential cannot be merely considered a minor perturbation to the free electron model Hamiltonian. Instead, it can result in a substantial effect on the electronic density of states, as shown in Refs. [10, 39].
@@ II.3 Clustering
In addition to this classification based on the static properties of the electron-lattice interaction, we here explore the dynamical aspects of this relationship by examining two distinct measures, namely the mean squared displacement (MSD) and inverse participation number (IPR). The spread of the wavepacket over time is measured by the MSD:
@@
a widely used method for analyzing scarred states or Anderson localized states in a disordered medium [40]. Here it is important to note that the measures discussed here deviate from their conventional definition by being determined as a function of time rather than time-independent, as typically seen in the studies of eigenfunctions. Furthermore, we can combine the time-evolving quantities into one two-dimensional times series, denoted as
@@
which paves the way to leveraging time-series ML methods for discerning various transport regimes.
Specifically, we apply $k$-means [41] clustering using dynamic time warping [42] to the mean-variance normalized set of series $\{\mathbf{F}(t)\}$ across different system variable settings. At its core, $k$-means is an algorithm to solve the optimization problem of partitioning a given set into $k$ clusters such that the in-cluster variance is minimized (for a detailed explanation of $k$-means and our method, we refer the reader to Appendix A). It is important to stress that this optimization is done in a fully unsupervised manner, meaning it only processes the raw time-series $\{\mathbf{F}(t)\}$ and is blind to the system variables used to generate the data. To ensure robustness and to mitigate the effect of statistical fluctuations, we average the clustering results of an ensemble of 10 time-series data sets, each generated using randomly initialized deformation potentials. This method allows us to objectively identify unique clusters corresponding to diverse dynamical regimes hidden within the Fröhlich Hamiltonian.

### v2

@@ Abstract
The intricate relationship between electrons and the crystal lattice is a linchpin in condensed matter, traditionally described by the Fröhlich model encompassing the lowest-order lattice-electron coupling. Recently developed quantum acoustics, emphasizing the wave nature of lattice vibrations, has enabled the exploration of previously uncharted territories of electron–lattice interaction not accessible with conventional tools such as perturbation theory. In this context, our agenda here is two-fold. First, we showcase the application of machine learning methods to categorize various interaction regimes within the subtle interplay of electrons and the dynamical lattice landscape. Second, we shed light on a nebulous region of electron dynamics identified by the machine learning approach and then attribute it to transient localization, where strong lattice vibrations result in a momentary Anderson prison for electronic wavepackets, which are later released by the evolution of the lattice. Overall, our research illuminates the spectrum of dynamics within the Fröhlich model, such as transient localization, which has been suggested as a pivotal factor contributing to the mysteries surrounding strange metals. Furthermore, this paves the way for utilizing time-dependent perspectives in machine learning techniques for designing materials with tailored electron–lattice properties.
@@ I Introduction
Anderson localization refers to the cessation of diffusive wave propagation in disordered systems [1]. On the historical front, Thouless theoretically posited [2] that at low temperatures, where inelastic processes are minimal, localization would result in higher resistance compared to that expected from ordinary elastic scattering. This insight later spurred the development of the scaling theory of Anderson localization for non-interacting electrons [3]. On the other hand, the conditions facilitating Anderson localization within an interacting system have been found to rely on several factors, including the strength of disorder, the dimensionality of the system [4], the range and type of interactions [5, 6, 7], and the time scales of the disorder potential dynamics [8, 9].
@@
The conundrum of whether systems localize or not was recognized early on by researchers like Gogolin [10, 11], Thouless [2], and also Anderson [1, 12]. For instance, the complex interplay between Anderson localization and lattice vibrations is observed in various random metal alloys and other disordered systems, such as crystalline organic semiconductors [13, 14] and halide perovskites [15]. The random fluctuations caused by lattice motion gradually disrupt the quantum interference necessary for electronic state localization, leading to what has been coined transient localization (for capturing the essential aspects, see, e.g., Ref. [9]). This phenomenon combines aspects of both Anderson localized and itinerant electron systems: Electronic transport is characterized by the successive cycles of localization and delocalization à la Anderson stemming from lattice vibrations that eventually result in reduced diffusion.
Whereas Anderson localization is typically explored within the framework of a tight-binding scheme featuring random on-site energies, the standard model for lattice vibrations is established by Fröhlich, which features linear coupling between an electron and the lattice. Conventionally, lattice vibrations are viewed through a number state perspective, but the coherent state representation introduced in Ref. [16], known as quantum acoustics, treats lattice vibrations as waves rather than individual phonons. This picture utilizing the coherent state basis is a valid way to treat the lattice vibrations fully quantum-mechanically, of equal, unassailable stature to the conventional Fock (number) state approach. A quantum lattice field in a number state has a well-defined amplitude, i.e., the number of quanta, but lacks knowledge of phase. On the other hand, the field defined by a coherent state has an equal amount of uncertainty in both amplitude and phase (a more detailed discussion on the coherent states can be found, e.g., in Refs. [17, 18, 19, 20]). However, even though these two pictures are equivalent at the most fundamental level, this duality is normally hidden by the approximations the two limits encourage. For example, a virtue of coherent states is that they are the closest quantum mechanical states to a classical description allowed by the uncertainty principle.
@@
The quantum-acoustical perspective unveils a duality between particle and wave pictures akin to quantum optics [17, 18, 19, 20]) established by Glauber [21]. Moreover, it allows for the electron–lattice interactions to be described in terms of a quasi-classical internal field, reminiscent of Bardeen and Shockley’s concept regarding dynamical lattice distortions in nonpolar semiconductors [22, 23]. In particular, the  deformation potential arising from lattice vibrations enables a quantum-coherent, nonperturbative treatment of charge carriers in coordinate space. In addition to recovering the results of the conventional Bloch-Grüneisen thory [16], the program of quantum acoustics has illuminated mysteries surrounding strange metals where transient localization plays a central role, such as T-linear resistivity at the Planckian limit surpassing the Mott–Ioffe–Regel threshold [24] and a shift in the Drude peak in the optical conductivity towards the infrared range [25]. Motivated by these advancements, we aim to identify various classes of dynamics hidden within the venerable Fröhlich model, which we express in the coherent state representation.
@@
The quantum acoustical approach above enables the generation of large amounts of time-dependent charge carrier wavefunctions as a function of the system parameters. Clustering, a common unsupervised learning technique, provides an effective means to explore the spectrum of carrier behavior by grouping similar dynamical profiles into clusters. In general, unsupervised machine learning (ML) methods have been established as a powerful tool to identify complex patterns in large unstructured data sets [26, 27, 28].
@@
In the broader landscape of ML applications in physics, our approach aligns with the recent uses of machine learning to understand and categorize complex physical phenomena, such as many-body localization and phase transitions [28, 29, 30, 31, 32, 33]. However, it is important to distinguish our work from the common narrative of “using ML to do physics”. Instead, our method uses ML as a tool that complements traditional analytical and numerical methods. This distinction underscores a shift from merely applying ML techniques to physics problems towards a more integrated approach where ML assists in how we conceptualize and explore physical systems.
To the authors’ knowledge, this study is the first to apply ML techniques for analyzing the dynamics of condensed matter systems through a time-dependent lens. Moreover, our approach not only goes beyond the established focus on eigenstates but also extends the application of ML to condensed matter systems outside of tight-binding models, such as spin chains.
@@
Our program is as follows. In Section II, we delineate the theoretical framework across three stages. We first put forward the concept of deformation potential (Section II.1), highlighting its significance as a palpable nonperturbative internal field for electrons (Section II.2). To facilitate the analysis of electron–lattice dynamics, we introduce a machine learning methodology in Section II.3. In Section III, we present our classification of wavepacket dynamics leveraging the ML approach, exploring variations in the strength of the electron–lattice interaction and illustrating a resulting “phase diagram”. Additionally, we conduct a detailed examination of one of the identified sectors connected to transient localization. Finally, we conclude our findings and discussions in Section IV.
@@ II Theory and Methods
where $c_{\mathbf{p}}$ $(c_{\mathbf{p}}^{\dagger})$ is the creation (annihilation) operator for electrons with momentum $\mathbf{p}$ and energy $\varepsilon_{\mathbf{p}}$ whereas $a_{\mathbf{q}}$ $(a_{\mathbf{q}}^{\dagger})$ is the creation (annihilation) operator for longitudinal acoustic phonons of wave vector $\mathbf{q}$ and energy $\hbar\omega_{\mathbf{q}}$, respectively. The  electron–phonon interaction is defined by its Fourier components $g_{\mathbf{q}}$. This Hamiltonian embodies the lattice $\mathbf{q}$, the electrons $\mathbf{p}$, and their lowest-order (linear) interaction that we next cast into the multimode coherent state basis of lattice degrees of freedom $|\chi\rangle$.
@@ II.1 Deformation Potential
The coherent state picture developed in Ref. [16] is the dual partner of the traditional number state description of electron–lattice dynamics. In this framework, each normal mode of lattice vibration with a wave vector $\mathbf{q}$ is associated with a coherent state $|\mathbf{q}\rangle$. At thermal equilibrium, each mode can be considered to be equilibrated with a heat bath at temperature $T$, giving thermal ensembles of coherent states where the average occupation of the mode $\langle n_{\mathbf{q}}\rangle_{\textrm{th}}$ is given by the Bose–Einstein distribution. Employing the independence of normal modes, entire lattice vibrations can be described as the product state of the coherent states of the normal modes, in other words, as a multimode coherent state $|\chi\rangle=\bigotimes_{\mathbf{q}}|\mathbf{q}\rangle$, as studied in Ref [34].
Even though the Fock state perspective focusing on the particle characteristics of lattice vibrations and the coherent state viewpoint emphasizing the wave nature are formally equivalent, the approximations they inspire are vastly different. For instance, a  common approach is to assume a direct product state $|\mathbf{p}\rangle\otimes|\chi\rangle$, combining the electronic state $|\mathbf{p}\rangle$ and the lattice state $|\chi\rangle$ while neglecting entanglement effects; this approach is equivalent to employing the time-dependent Hartree approximation. Moreover, we only consider the longitudinal acoustic branch of lattice vibrations. Then, as  detailed in Ref. [16], the quasi-classical limit of quantum acoustics unveils a real-space, time-dependent description of electron–lattice interaction in terms of the deformation potential.
@@
where $\varphi_{\mathbf{q}}$ is the phase of the coherent state $|\mathbf{q}\rangle$. Furthermore, we assume the phases $\varphi_{\mathbf{q}}$ to be uniformly distributed random variables and employ the Debye model, assuming the linear dispersion $\omega_{\mathbf{q}}=v_{s}|\mathbf{q}|$, where $v_{s}$ is the speed of sound. Therefore, the time dependence of the deformation potential is governed by the following wave equation:
@@
The deformation potential is a peculiar object. It is homogeneously random, meaning that the probability distribution of potential values $V_{D}$ does not depend on a position $\mathbf{r}$ or time $t$ (with the assumption of random phases). Therefore, each spatio-temporal patch of the potential is statistically indistinguishable from another. The typical length scale of the spatial correlations is determined by its largest wavenumber components $\sim q_{D}$. Similarly, the typical timescale of the potential change is determined by its largest frequency components $\sim\omega_{D}$. This special type of spatial-temporal correlation sets the deformation potential apart from other types of lattice distortions, as commonly investigated in the context of Anderson localization [1].
@@ II.2 Electron Dynamics
where $m^{*}$ is the effective (band) mass of the electron and $V_{D}(\mathbf{r},t)$ is the deformation potential given by Equation (2). This effective Hamiltonian $\mathcal{H}_{0}$ represents the electron component of the Fröhlich model defined previously in Equation (1) within the framework of the effective mass approximation.
Our investigation of electron dynamics under the defined effective Hamilton $\mathcal{H}_{0}$ approaches the issue from the point of view of Gaussian wavepackets that are a common tool for analyzing time-dependent aspects of a quantum system [38, 39], for instance in the studies of quantum optics [18, 17], scarring [40, 41, 42], and  branched flow [43, 44, 45]. Here, we choose the following test Gaussian for representing the charge carrier:
@@
where $\mathcal{N}$ is the normalization factor, and $\bm{\sigma}=(\sigma_{x}^{-1},\sigma_{y}^{-1})$ describes the initial width of the wavepacket. Without  loss of generality, we can choose to launch the test wavepacket into the x direction with the Fermi momentum, thus $\mathbf{k}=(k_{F},0)$, where $k_{F}$ is the Fermi wavevector. The memory of the initial form of the wavepacket is quickly lost in the chaotic potential and its exact form is unimportant.
@@
Figure 2 illustrates the charge carrier wave, originally a Gaussian as described in Equation (4), evolving under the influence of the dynamic lattice wave field, which converts the always-accessible wave nature of lattice vibrations into something valuable, a point where the quantum-acoustical perspective becomes tangible. Within this kind of Wave-on-Wave (WoW) approach, as detailed in Refs. [25, 24], one winds up solving two interacting equations of motions simultaneously: one for the lattice and one for the electron (i.e., the time-dependent Schrödinger equation for the charge carrier and the wave equation for the lattice vibrations).
Figure 2: The quantum acoustical Wave-on-Wave (WoW) approach to charge carrier dynamics. An electron wavepacket propagates atop a deformation potential, which itself evolves according to the wave equation. As it traverses this shifting acoustic landscape shaped by acoustic deformations, the electron undergoes quasi-elastic scattering akin to impurity scattering.
@@
On the other hand, the comparison of the kinetic energy of the electron (a fair approximation is given by the Fermi energy) with the root mean square of the deformation potential determines whether the lattice vibration and its resultant electron scattering can be treated perturbatively or not. In essence, the deformation potential cannot be merely considered a minor perturbation to the free-electron model Hamiltonian. Instead, it can result in a substantial effect on the electronic density of states, as shown in Refs. [16, 48].
@@ II.3 Clustering
In addition to this classification based on the static properties of the electron–lattice interaction, we here explore the dynamical aspects of this relationship by examining two distinct measures, namely the mean squared displacement (MSD) and inverse participation ratio (IPR). The spread of the wavepacket over time is measured by the MSD:
@@
a widely used method for analyzing scarred states or Anderson localized states in a disordered medium [49]. Here, it is important to note that the measures discussed here deviate from their conventional definition by being determined as a function of time rather than time-independent, as typically seen in the studies of eigenfunctions. Furthermore, we can combine the time-evolving quantities into one two-dimensional times series, denoted as
@@
which paves the way to leveraging ML methods for time series to discern various transport regimes.
[…]